OpenStudy (dmezzullo):

Solve: 6(x + 4) = 8(x − 2) −4 4 −20 20

OpenStudy (dmezzullo):

@uri @mikala1

OpenStudy (anonymous):

6x+24 = 8x-16 40= 2x x= 40/2 x=20

OpenStudy (dmezzullo):

tyvm